OpenStudy (anonymous):

Jen bikes the 36 km to the resort averaging a certain speed. The return trip is made at a speed 3 khp slower. The total time for the round trip is 7 hours. Find Jen's Average speed on each part of the trip.

OpenStudy (rushwr):

@imqwerty @ChiefArnav

OpenStudy (arindameducationusc):

use speed=distance/time

OpenStudy (arindameducationusc):

here distance= 36+36

OpenStudy (arindameducationusc):

time=7hours

OpenStudy (arindameducationusc):

and speed is x+(x-3)

OpenStudy (anonymous):

How do you get the speed?

OpenStudy (arindameducationusc):

72/7=x+(x+3). This should be the way........

OpenStudy (arindameducationusc):

find x which is the answer

OpenStudy (anonymous):

Thanks. :)